OpenStudy (anonymous):

HELP ME GUYS PLEASE, I'M BEGGING YOU HUHU :( 1. A right circular cylinder has 4in as its diameter and 6in as its height. a. Find the lateral area b. Find the volume of the regular hexagonal prism inscribed inside the cylinder 2. A rectangular solid is 4ft long and 3ft wide a. Find the volume if its diagonal is square root of 41.

OpenStudy (anonymous):

24pi in^2 is the lateral area

OpenStudy (anonymous):

how about letter b?

OpenStudy (anonymous):

b. \[lateral~area=2 \pi~r~h=2*\pi 2*6=24\pi~ inch ^{2}\] 

OpenStudy (anonymous):

 volume of cuboid=l*b*h=4*3*4 =?